The point estimate for the proportion of 16-kbit dynamics RAMs that function correctly for at least 1000 hours based on a sample size 100 is .91. Find a 95% confidence interval on p.


1
Expert's answer
2022-05-02T03:26:03-0400

The critical value for "\alpha = 0.05" is "z_c = z_{1-\alpha/2} = 1.96."

The corresponding confidence interval is computed as shown below:


"CI(Proportion)=(\hat{p}-z_c\times\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}},"

"\hat{p}+z_c\times\sqrt{\dfrac{\hat{p}(1-\hat{p})}{n}})"

"=(0.91-1.96\times\sqrt{\dfrac{0.91(1-0.91)}{100}},"

"0.91+1.96\times\sqrt{\dfrac{0.91(1-0.91)}{100}})"

"=(0.854,0.966)"

Therefore, based on the data provided, the 95% confidence interval for the population proportion is "0.854 < p < 0.966," which indicates that we are 95% confident that the true population proportion "p" is contained by the interval "(0.854, 0.966)."